Trevor always begins the day with a strawberry milkshake (milk (x1 ) and strawberries(x2) mixed in proportion 1:5). His income is equal to m=200, and one strawberry costs p2=1. Suppose the price of milk drops from p1=15 to p1=5. We are going to decompose the total effect into substitution effect and income effect.

a) What is the total change in demand for milk?

b) What is the substitution effect?

c) What is the income effect?
Answer #1

Given that the proportion of milk and strawberry for milk shake is 1:5.

If quantity of milk is X1 then quantity of strawberry (X2) will be 5×X1, i.e., X2=5X1... Equation 1

And also that the milk and strawberry are complementary goods as milk is of no use without strawberry and vice versa.

Budget equation will be as follows:

P1×X1+P2×5X1=M.... Equation 2

Given M=200, P1=15 & P2=1

Putting values in Equation 2

15×X1+1×5X1=200

X1=10 & X2=50(from equation 1)

Answer a)

With change in P1 from 15 to 5

Again putting values in Equation 2

5×X1+1×5X1=200

X1=20 & X2=100.

Total change in demand of milk is increase from 10 units to 20 units.

Answer b)

Milk and strawberry are complementary goods here therefore there would be no substitution effect.

Answer c)

Since there is no substitution effect, total effect is equal to income effect.